OpenStudy (anonymous):

Could I find the missing side lengths of a right triangle if I only know one side length and one angle measure?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

But the angle measure has to be one of the other angles, not the right angle

OpenStudy (anonymous):

Can you gve me an example please?

OpenStudy (mathmate):

 If \( \theta = 25\) degrees, and c=20, then \( b=c\ sin(\theta) \) and \( a=c\ cos(\theta) \).

OpenStudy (anonymous):

im so confused with these ratios.

OpenStudy (mathmate):

Are these sin() cos() ratios new to you, or you have not done them yet?

OpenStudy (anonymous):

i just dont understand them. i know what they are, but how do i use them????

OpenStudy (mathmate):

The standard trigonometric ratios, referring to the figure: \( sin(\theta)=opposite/hypotenuse = b/c \) \( cos(\theta)=adjacent/hypotenuse = a/c \) \( tan(\theta)=opposite/adjacent = a/c \) Draw a diagram to which you can refer when you do these exercises. They are not easy to memorize, but when you will have done some exercises, they will sit in your memory automatically. Refer to your diagram often. Do not force yourself to memorize it because it will actually slow you down when you with them later on.

OpenStudy (mathmate):

erratum: tan(θ)=opposite/adjacent=\(b/a\)

OpenStudy (anonymous):

o i just need to know how to apply the sin cos and tan ratios. Possibly the recipercals also.

OpenStudy (phi):

Here is Khan's intro http://www.khanacademy.org/video/basic-trigonometry?playlist=Trigonometry (and the video following this one) and here is how to use trig functions http://www.khanacademy.org/video/using-trig-functions?playlist=Trigonometry

OpenStudy (mathmate):

The videos are excellent source of knowledge. You will find that it's worthwhile investing a few minutes on them.

OpenStudy (anonymous):

ok i can not veiw the in school, but i will soon thnks!

OpenStudy (mathmate):

 So here are the six ratios, including the reciprocals. Note that I have put the length of the hypotenuse as 1. If the hypotenuse were L, then the opposite side would be L sin(\(\theta\)) and the adjacent side would be L cos(\(\theta\)).